OpenStudy (peace):

How to solve abs(e^(2x) - (1/x+2)) = 2

OpenStudy (anonymous):

\[| e ^{2x} - 1/(x+2)| = 2\] absolute value mean that it's =( 2 ) or =( -2 ) so, solve for both and you'll have yur answer. got it ?

OpenStudy (peace):

yea i'm trying to do that, but i'm not really getting anywhere

OpenStudy (anonymous):

Ok. So, for the first equality: \[e ^{2x} - 1/(x+2) = 2\] you can take out the fraction by inverting everything:\[e^{-2x} - (x+2) = 1/2\]\[e^{-2x} - x = 5/2\] to eliminate the {e}, multiply everything by Natural Log (ln)\[-2x - \ln x = \ln (5/2)\] solve for x and there might be more than 1 solution

OpenStudy (peace):

how do i go on from -2x = ln (5x/2)

OpenStudy (anonymous):

sorry, i can't remember/find how to solve :/

OpenStudy (peace):

aw it's ok, thanks for your help though :)

OpenStudy (anonymous):

np, if i find somewhere i'll post it here (: